# On a combinatorial classification of fine compactified universal Jacobians

### v1

@@ Abstract.
Extending the definition of $V$-stability conditions, given by Viviani in the recent preprint [Viv23], we introduce the notion of universal stability conditions. Building on results by Pagani and Tommasi in [PT23], we show that fine compactified universal Jacobians, that is, fine compactified Jacobians over the moduli spaces of stable pointed curves ¯$\overline{\mathcal{M}}_{g,n}$, are combinatorially classified by universal stability conditions.
@@ 1. Introduction
In this paper, we extensively use results from the recent preprints [PT23] and [Viv23], by Pagani-Tommasi and Viviani respectively, to give a complete classification of fine compactified universal (over each moduli stack ¯$\overline{\mathcal{M}}_{g,n}$ of stable $n$-pointed curves of genus $g$) Jacobians in terms of universal $V$-stability conditions. This addresses [Viv23, Open Question (4), p.7]. In doing so, we answer, for all $g\geq 2$, the following (raised in [PT23, Section $10$]):
@@
Compactified Jacobians have been constructed via different methods by Oda-Seshadri [OS79], for a single nodal curve, by Altman-Kleiman [AK80], Esteves [Est97], Simpson [Sim94] for families of curves over a scheme, and by Caporaso [Cap94], Pandharipande [Pan95] for the universal family. Following [Viv23], we will call these fine classical compactified Jacobians (see Definitions 3.6 and 2.14 for a precise definition).
The first construction of space of stability conditions for fine classical compactified universal Jacobians was given in [KP19]. Each of the maximal dimensional chambers of the stability space defined in loc.cit. corresponds to a fine compactified universal Jacobian, that we will refer to as fine classical compactified universal Jacobians (Definition 3.6). (These were earlier constructed in [MRV15] and [KP17]). The existence of non-classical fine compactified universal Jacobians was first observed in [PT22], and the examples were given when $g=1$ and $n\geq 6$. However, to the best of our knowledge, in all known examples of non-classical fine compactified universal Jacobians, the fibres over each geometric point of ¯$\overline{\mathcal{M}}_{g,n}$ are classical (in the sense of the previous paragraph).
Recently, Pagani and Tommasi gave in [PT23] a complete combinatorial classification of all fine compactified universal Jacobians. This classification is fairly complicated as it uses a large amount of data. For example trying to approach Question 1 by means of this combinatorial description seems very hard.
Inspired by the more recent paper of Viviani [Viv23], in this work we introduce the notion of universal stability conditions (Definition 2.4), whose restriction to a curve $X$ (a $c$-stability condition, Definition 2.8) agrees with the notion of a $V$-stability condition introduced in [Viv23]. We show that the set of universal stability conditions is precisely the combinatorial counterpart of the set of fine compactified universal Jacobians. It is by means of this explicit combinatorial description that we are able to fully answer Question $\star$1.
@@
Viceversa, let $\overline{\mathcal{J}}_{g,n}$ be a degree $d$ fine compactified universal Jacobian. We can obtain its associated universal stability condition $\mathfrak{m}(\overline{\mathcal{J}}_{g,n})$ (Definition 3.2) by taking the minimum degree, among the sheaves $[\mathcal{F}]\in\overline{\mathcal{J}}_{g,n}$, of their restrictions to the irreducible components of the fibres over the vine curves (i.e. curves with two irreducible components and no self intersections).
@@
Similarly to the case of $V$-stability conditions on a graph $\Gamma$, a $G_{g,n}$-morphisms compatible classical stability condition $\phi$ induces a $c$-stability condition $\mathfrak{m}^{\Gamma}(\phi)$ (Definition 2.16).
However, not all universal stability conditions arise as limits of classically induced $c$-stability conditions. In fact, in Section 3.4, we find suitable non classical $c$-stability conditions on trivalent graphs in $G_{g,n}$, which are restriction of a universal stability condition. Furthermore, we show that, given any genus $g$, such graphs have the minimal number of markings $n$ for this to be possible. Hence, we deduce our main result, answering Question $\star$1:
@@
    $\begin{Bmatrix}\textup{Degree $d$ fine compactified universal}\\ \textup{Jacobians of type $(g,n)$ whose fibres}\\ \textup{over each pointed curve is classical}\end{Bmatrix}\hookrightarrow\begin{Bmatrix}\textup{Degree $d$ fine compactified}\\ \textup{universal Jacobians of type $(g,n)$}\end{Bmatrix}$
@@
    • $n=0$ (shown in [PT23, Theorem 9.6], see also Example 3.17).
    • $g=1$ (shown in [PT22, Proposition 3.15]).
@@
We can also ask the related question of when the set of fine classical compactified universal Jacobians is a proper subset of the set of all fine compactified universal Jacobians. The $n$ given above is sharp for the latter inclusion to be strict whenever $g\geq 3$. Indeed, by Proposition 3.15, every fine compactified universal Jacobian over ¯$\overline{\mathcal{M}}_{3,1}$ is classical. In Corollary 3.14 we show that for $g=2$ the latter inclusion is strict also when $n=4,5$. This, combined with Theorem B, completely answers the related question.
@@ 1.1. Acknowledgements
I thank my supervisor Nicola Pagani for the many useful insights and conversations. I would also like to acknowledge Rhys Wells for his contribution on the computational side.
@@ 2.1. Graphs
We recall the notation concerning graphs from [MMUV22] and [Viv23], which we will adopt throughout the paper.
@@ 2.2. Universal and c-stability conditions
In this section we introduce the notion of universal stability condition on $G_{g,n}$ and restrict it to a $c$-stability condition on a graph $\Gamma$. We show that the notion is equivalent to the one of $V$-stability condition on the graph $\Gamma$ that is compatible with morphisms of $G_{g,n}$. Moreover, a classical $\phi$-stability condition compatible with $G_{g,n}$-morphisms on such a graph always induces a $c$-stability condition, while viceversa is not necessarily true.
@@
Let $\{\mathfrak{m}^{\Gamma}\}_{\Gamma\in G_{g,n}}$ be an assignment of $c$-stability conditions such that, for each $\Gamma,\Gamma^{\prime}\in G_{g,n}$ and $(e,h,A)\in\mathcal{D}(\Gamma)\cap\mathcal{D}(\Gamma^{\prime})$, $\mathfrak{m}_{(e,h,A)}^{\Gamma}=\mathfrak{m}_{(e,h,A)}^{\Gamma^{\prime}}$. Taking the limit over $G_{g,n}$ gives a universal stability condition $\mathfrak{m}$, which coincides with $\mathfrak{m}^{\Gamma}$ on each $\Gamma\in G_{g,n}$. That is, the following holds
@@
We, now, want to show what is the relation between $V$-stability conditions, as defined in [Viv23, Definition 1.4], and $c$-stability conditions on a given graph $\Gamma\in G_{g,n}$.
@@
We recall that morphisms in $G_{g,n}$ are contractions of prescribed edges followed by an isomorphism, thus they induce bijections on the level of subsets of vertices. That is, if $c:\Gamma\to\Gamma^{\prime}$ is a morphism, for each $Z\subset V(\Gamma^{\prime})$, there exists only a $W\subset V(\Gamma)$ such that $c(W)=Z$. This fact ensures the following definition is well posed.
@@
First, we show that for each degree $d$ $c$-stability condition $\mathfrak{m}^{\Gamma}$, $\beta(\mathfrak{m}^{\Gamma})$ is indeed a $V$-stability condition of degree $d$ on $\Gamma$ compatible with $G_{g,n}$-morphisms. As the compatibility requirement is trivially satisfied by definition, we only need to prove that $\beta(\mathfrak{m}^{\Gamma})=\{\mathfrak{m}^{\Gamma}_{(\textup{val}(Z),g(Z),A(Z))}:Z\in\mathcal{C}(\Gamma)\}$ satisfies Properties (1) and (2) of [Viv23, Definition 1.4].
@@
Notice that, in general, given a graph $\Gamma\in G_{g,n}$, not all the $V$-stability conditions on $\Gamma$ are necessarily compatible with $G_{g,n}$-morphisms. An example is given by [Viv23, Example 1.27].
As previously announced, we now show how $G_{g,n}$-morphisms compatible classical stability conditions induce $c$-stabilities, on a given graph $\Gamma$ as well as in the universal case. Recalling the language of [OS79], we give the following:
@@
A degree $d$ classical stability condition $\phi^{\Gamma}$ on a graph $\Gamma\in G_{g,n}$ is a real divisor $\phi^{\Gamma}\in\textup{Div}(\Gamma)\otimes\mathbb{R}=\mathbb{R}^{V(\Gamma)}$ of total degree $|\phi^{\Gamma}|=\sum_{v\in V(\Gamma)}\phi^{\Gamma}_{v}=d$.
Such a stability condition is said to be general if it further satisfies the property
@@
Let $\{\phi^{\Gamma}\}_{\Gamma\in G_{g,n}}$ be a family of classical stability conditions which are compatible with contractions and automorphisms in the sense of [KP19, Section $3$]. A degree $d$ universal classical stability condition is the limit
@@
Let $\Gamma\in G_{g,n}$ be a graph, and let $\phi$ be a degree $d$ classical general stability condition on $\Gamma$, compatible with $G_{g,n}$-morphisms. We define the $\phi$-induced $c$-stability condition $\mathfrak{m}^{\Gamma}(\phi)$ on $\Gamma$ by the assignment
$\mathfrak{m}^{\Gamma}(\phi):=\{\mathfrak{m}_{(e,h,A)}^{\Gamma}(\phi):=\Bigl{\lceil}\phi_{Z(e,h,A)}-\frac{e}{2}\Bigr{\rceil}:(e,h,A)\in\mathcal{D}(\Gamma)\},$
@@
The induced stability condition $\mathfrak{m}^{\Gamma}(\phi)$ is the image, via the map of Proposition 2.12, of the $V$-stability condition
$\mathfrak{n}^{\Gamma}(\phi):=\{\mathfrak{n}^{\Gamma}(\phi)_{Z}=\Bigl{\lceil}\phi^{\Gamma}_{Z}-\frac{\textup{val}(Z)}{2}\Bigr{\rceil}:Z\in\mathcal{C}(\Gamma)\},$
induced by $\phi$, defined in [Viv23, Example 1.5].
Thus, to show that $\mathfrak{m}^{\Gamma}(\phi)$ is well defined, it is sufficient to notice that whenever $\phi$ is compatible with $G_{g,n}$-morphisms, so is the induced $V$-stability condition. Indeed, this is true by the following observation. Let $\Gamma^{\prime}\in G_{g,n}$ and $c,\tilde{c}:\Gamma\to\Gamma^{\prime}$ be morphisms. For each $Z\in V(\Gamma^{\prime})$,
@@
A more explicit form for the induced stability condition $\mathfrak{m}^{\Gamma}(\phi)$ will prove useful for constructing examples in the following section, thus we show:
@@
Let $g\geq 2$, and let $\Gamma\in G_{g,n}$ be a $2$-connected graph (i.e. a graph without separating edges). Let $\phi$ be a classical general degree $d$ stability condition on $\Gamma$, compatible with $G_{g,n}$-morphisms. Then, the $\phi$-induced $c$-stability condition $\mathfrak{m}^{\Gamma}(\phi)$ satisfies
$\mathfrak{m}_{(e,h,A)}^{\Gamma}(\phi)=\Bigl{\lceil}\frac{(e+2h-2)(d+1-g)+(2g-e-2h)\phi_{A}-(e+2h-2)\phi_{A^{\mathsf{c}}}}{2g-2}\Bigr{\rceil}+h-1,$
where $\phi_{A}=\sum_{v:A(v)\subset A}\phi_{v}$, for $v\in V(\Gamma)$.
@@
We observe that the $V$-stability condition $\mathfrak{n}(\phi)$ induced by a classical stability condition $\phi$ of [Viv23, Example 1.5] depends on the stability polytope containing $\phi$, rather than on the specific $\phi$.
Let $c:\Gamma^{\prime}\to\Gamma$, be a morphism, and let $\phi$ be a classical stability condition on $\Gamma$. By generality of $\phi$, there exists a stability condition $\phi^{\prime}$ on $\Gamma^{\prime}$ such that the pushforward $c(\phi^{\prime})$ is in the same polytope as $\phi$. Therefore it is enough to prove our statement on the initial objects of $G_{g,n}$, that is, trivalent graphs.
So, we assume that $\Gamma$ is trivalent. By compatibility with $G_{g,n}$-morphisms, since $\Gamma$ is $2$-connected, we have $\phi_{v_{1}}=\phi_{v_{2}}$ for any pair of unlabeled genus 00 vertices $v_{1},v_{2}$. Hence, the $c$-stability condition induced by $\phi$ on a $2$-connected trivalent graph in $G_{g,n}$ has elements
$\mathfrak{m}_{(e,h,A)}^{\Gamma}(\phi)=\Bigl{\lceil}\frac{k(d+1-g)+(2g-2-k)\phi_{A}-k\phi_{A^{\mathsf{c}}}}{2g-2}\Bigr{\rceil}+h-1,$
@@
Let $\phi=\{\phi^{\Gamma}\}_{\Gamma\in G_{g,n}}$ be a collection of a $G_{g,n}$-morphisms compatible classical stability condition $\phi^{\Gamma}$ for each $\Gamma$ in $G_{g,n}$. If, for each pair of graphs $\Gamma,\Gamma^{\prime}$, we have
@@
for any $(e,h,A)\in\mathcal{D}(\Gamma)\cap\mathcal{D}(\Gamma^{\prime})$, we define the $\phi$-induced universal stability condition $\mathfrak{m}(\phi)$ as
@@
Clearly, not only each element $\phi^{\Gamma}$ of a contractions and automorphisms compatible collection $\{\phi^{\Gamma}\}_{\Gamma\in G_{g,n}}$ is compatible with $G_{g,n}$-morphisms, but their collection induces a family $\{\mathfrak{m}^{\Gamma}\}_{\Gamma\in G_{g,n}}$, whose integral data $\mathfrak{m}_{(e,h,A)}^{\Gamma}$ is the same for every $\Gamma$ such that $(e,h,A)\in\mathcal{D}(\Gamma)$. As a consequence, we obtain that every classical universal stability condition, as defined by [KP19, Section 3], induces a universal stability condition.
Observe, however, that not only universal classical stability conditions induce universal stability conditions. In fact, Example 6.15 in [PT22] shows that there exists universal stability conditions non induced by a universal classical stability condition, whose restriction to any graph is induced by a classical stability condition (see also Example 3.12 below).
The following lemma allows us to check whether a $c$-stability condition on a graph $\Gamma$ is classical only on half of the elements of $\mathcal{D}(\Gamma)$.
@@
Let $\mathfrak{m}^{\Gamma}$ be a degree $d$ $c$-stability condition on a $2$-connected graph $\Gamma\in G_{g,n}$, with $g\geq 2$, and let $(e,h,A)\in\mathcal{D}(\Gamma)$. Let $\phi\in\mathbb{R}^{n}$ be a classical stability condition on $\Gamma$ such that $\mathfrak{m}_{(e,h,A)}^{\Gamma}=\mathfrak{m}_{(e,h,A)}^{\Gamma}(\phi)$. Then $\mathfrak{m}^{\Gamma}_{(e,g+1-e-h,A^{\mathsf{c}})}=\mathfrak{m}^{\Gamma}_{(e,g+1-e-h,A^{\mathsf{c}})}(\phi)$.

### v2

@@ Abstract.
Extending the definition of $V$-stability conditions, given by Viviani in the recent preprint [Viv], we introduce the notion of universal stability conditions. Building on results by Pagani and Tommasi in [PT], we show that fine compactified universal Jacobians, that is, fine compactified Jacobians over the moduli spaces of stable pointed curves ¯$\overline{\mathcal{M}}_{g,n}$, are combinatorially classified by universal stability conditions.
@@ 1. Introduction
In this paper, we extensively use results from the recent preprints [PT] and [Viv], by Pagani-Tommasi and Viviani respectively, to give a complete classification of fine compactified universal (over each moduli stack ¯$\overline{\mathcal{M}}_{g,n}$ of stable $n$-pointed curves of genus $g$) Jacobians in terms of universal $V$-stability conditions. This addresses [Viv, Open Question (4), p.7]. In doing so, we answer, for all $g\geq 2$, the following (raised in [PT, Section $10$]):
@@
Compactified Jacobians have been constructed via different methods by Oda-Seshadri [OS79], for a single nodal curve, by Altman-Kleiman [AK80], Esteves [Est01], Simpson [Sim94] for families of curves over a scheme, and by Caporaso [Cap94], Pandharipande [Pan95] for the universal family. Following [Viv], we will call these classical fine compactified Jacobians (see Definitions 3.6 and 2.142.16 for a precise definition).
The first construction of space of stability conditions for classical fine compactified universal Jacobians was given in [KP19]. Each of the maximal dimensional chambers of the stability space defined in loc.cit. corresponds to a fine compactified universal Jacobian, that we will refer to as classical fine compactified universal Jacobians (Definition 3.6). (These were earlier constructed in [Mel16] and [KP17]). The existence of non-classical fine compactified universal Jacobians was first observed in [PT23], and the examples were given when $g=1$ and $n\geq 6$. However, to the best of our knowledge, all known examples of fine compactified universal Jacobians have fibres over each geometric point of ¯$\overline{\mathcal{M}}_{g,n}$ that are classical compactified Jacobians (in the sense of the previous paragraph). We call such compactified universal Jacobians semi-classical (Definition 3.6).
Recently, Pagani and Tommasi gave in [PT] a complete combinatorial classification of all fine compactified universal Jacobians. This classification is fairly complicated as it uses a large amount of data. For example trying to approach Question $\star$1 by means of this combinatorial description seems very hard.
Inspired by the more recent paper of Viviani [Viv], in this work we introduce the notion of universal stability conditions (Definition 2.4), whose restriction to a curve $X$ (a $c$-stability condition, Definition 2.8) agrees with the notion of a $V$-stability condition introduced in [Viv]. We show that the set of universal stability conditions is precisely the combinatorial counterpart of the set of fine compactified universal Jacobians. It is by means of this explicit combinatorial description that we are able to fully answer Question $\star$1.
@@
Viceversa, let $\overline{\mathcal{J}}_{g,n}$ be a degree $d$ fine compactified universal Jacobian. We can obtain its associated universal stability condition $\mathfrak{m}(\overline{\mathcal{J}}_{g,n})$ (Definition 3.2) by taking, for every components of all the marked vine curves (i.e. curves with two irreducible components and no self intersections) $[X,\{p_{1},\ldots,p_{n}\}]\in\overline{\mathcal{M}}_{g,n}$, the minimum degrees of sheaves $F$ such that $[X,\{p_{1},\ldots,p_{n}\},F]\in\overline{\mathcal{J}}_{g,n}$.
@@
Similarly to the case of $V$-stability conditions on a graph $\Gamma$, a $G_{g,n}$-morphisms compatible numerical polarisation $\phi^{\Gamma}$ induces a classical $c$-stability condition $\mathfrak{m}^{\Gamma}(\phi^{\Gamma})$ (Definition 2.16).
However, not all universal stability conditions arise as limits of classical $c$-stability conditions. In fact, in Section 3.4, we find suitable non classical $c$-stability conditions on trivalent graphs in $G_{g,n}$, which are restriction of a universal stability condition. Furthermore, we show that, given any genus $g$, such graphs have the minimal number of markings $n$ for this to be possible. Hence, we deduce our main result, answering Question $\star$1:
@@
    $\begin{Bmatrix}\textup{Degree $d$ semi-classical fine compactified}\\ \textup{universal Jacobians of type $(g,n)$}\end{Bmatrix}\hookrightarrow\begin{Bmatrix}\textup{Degree $d$ fine compactified}\\ \textup{universal Jacobians of type $(g,n)$}\end{Bmatrix}$
@@
    • $n=0$ (shown in [PT, Theorem 9.6], see also Example 3.17).
    • $g=1$ (shown in [PT23, Proposition 3.15]).
@@
More specifically, given any fine compactified universal Jacobian , we can decide whether it is classical or semi-classical by analysing the system of inequalities induced by its associated universal stability condition, via Proposition 2.17.
We can also ask the related question: what are the pairs $(g,n)$ such that the set of classical fine compactified universal Jacobians is strictly included in the set of all fine compactified universal Jacobians? (The answer to this was already known for $g\leq 1$ by work of [PT23] and for $n=0$ by work of [PT]).
The $n$ given in Theorem B is sharp for the latter inclusion to be strict whenever $g\geq 3$. Indeed, by Proposition 3.15, every fine compactified universal Jacobian over ¯$\overline{\mathcal{M}}_{3,1}$ is classical. In Corollary 3.14 we show that for $g=2$ the latter inclusion is strict also when $n=4,5$. This, combined with Theorem B, completely answers the related question.
@@ 1.1. Acknowledgements
I thank my supervisor Nicola Pagani for the many useful insights and conversations. I would like to acknowledge Rhys Wells for his contribution on the computational side. I am also grateful to Filippo Viviani for all the helpful comments and suggestions on the previous version of the paper.
@@ 2.1. Graphs
We recall the notation concerning graphs from [MMUV22] and [Viv], which we will adopt throughout the paper.
@@ 2.2. Universal and c-stability conditions
In this section we introduce the notion of universal stability condition on $G_{g,n}$ and restrict it to a $c$-stability condition on a graph $\Gamma$. We show that the notion is equivalent to the one of $V$-stability condition on the graph $\Gamma$ that is compatible with morphisms of $G_{g,n}$. Moreover, a classical numerical polarisation compatible with $G_{g,n}$-morphisms on such a graph always induces a $c$-stability condition, while viceversa is not necessarily true.
@@
Let $\{\mathfrak{m}^{\Gamma}\}_{\Gamma\in G_{g,n}}$ be an assignment of $c$-stability conditions that agree on common domains, i.e. such that, for each $\Gamma,\Gamma^{\prime}\in G_{g,n}$ and $(e,h,A)\in\mathcal{D}(\Gamma)\cap\mathcal{D}(\Gamma^{\prime})$, $\mathfrak{m}_{(e,h,A)}^{\Gamma}=\mathfrak{m}_{(e,h,A)}^{\Gamma^{\prime}}$. Taking the limit over $G_{g,n}$ gives a universal stability condition $\mathfrak{m}$, which coincides with $\mathfrak{m}^{\Gamma}$ on each $\Gamma\in G_{g,n}$. That is, the following holds
@@
We, now, want to show what is the relation between $V$-stability conditions, as defined in [Viv, Definition 1.4], and $c$-stability conditions on a given graph $\Gamma\in G_{g,n}$.
@@
We recall that morphisms in $G_{g,n}$ are contractions of prescribed edges followed by an isomorphism, thus they induce surjections on the level of subsets of vertices. In particular, if $c:\Gamma\to\Gamma^{\prime}$ is a morphism, for each $Z\subset V(\Gamma^{\prime})$, there exists a maximal $W\subset V(\Gamma)$ such that $c(W)=Z$. This fact ensures the following definition is well posed.
@@
First, we show that for each degree $d$ $c$-stability condition $\mathfrak{m}^{\Gamma}$, $\beta(\mathfrak{m}^{\Gamma})$ is indeed a $V$-stability condition of degree $d$ on $\Gamma$ compatible with $G_{g,n}$-morphisms. As the compatibility requirement is trivially satisfied by definition, we only need to prove that $\beta(\mathfrak{m}^{\Gamma})=\{\mathfrak{m}^{\Gamma}_{(\textup{val}(Z),g(Z),A(Z))}:Z\in\mathcal{C}(\Gamma)\}$ satisfies Properties (1) and (2) of [Viv, Definition 1.4].
@@
Notice that, in general, given a graph $\Gamma\in G_{g,n}$, not all the $V$-stability conditions on $\Gamma$ are necessarily compatible with $G_{g,n}$-morphisms, as shown by [Viv, Example 1.27]. Indeed, let $\Gamma$ be the graph in loc. cit. decorated with markings $\{1\}$ and $\{2\}$ at the vertices $5$ and $6$ respectively, and let $\mathfrak{n}^{\Gamma}$ be the $V$-stability condition associated to the mildly superadditive function $\beta$. Notice that $\Gamma$ is isomorphic to a graph in $G_{3,2}$. Consider the morphisms $c:\Gamma\to V(3,0,\{2\})$, that contracts edges $\{\overline{16},\overline{13},\overline{24},\overline{45}\}$ and $\tilde{c}:\Gamma\to V(3,0,\{2\})$, that contracts edges $\{\overline{16},\overline{12},\overline{35},\overline{45}\}$, in the notation of loc. cit. Then, we observe that
$10=\mathfrak{n}^{\Gamma}_{c^{-1}(\{v_{1}\})}\neq\mathfrak{n}^{\Gamma}_{\tilde{c}^{-1}(\{v_{1}\})}=$ 1 .
As previously announced, we now show how $G_{g,n}$-morphisms compatible numerical polarisations induce $c$-stabilities, on a given graph $\Gamma$ as well as in the universal case. Recalling the language of [OS79], we give the following:
@@
A degree $d$ numerical polarisation $\phi^{\Gamma}$ on a graph $\Gamma\in G_{g,n}$ is a real divisor $\phi^{\Gamma}\in\textup{Div}(\Gamma)\otimes\mathbb{R}=\mathbb{R}^{V(\Gamma)}$ of total degree $|\phi^{\Gamma}|=\sum_{v\in V(\Gamma)}\phi^{\Gamma}_{v}=d$.
Such a polarisation is said to be general if it further satisfies the property
@@
Let $\{\phi^{\Gamma}\}_{\Gamma\in G_{g,n}}$ be a family of numerical polarisations which are compatible with contractions and automorphisms in the sense of [KP19, Section $3$]. A degree $d$ universal numerical polarisation is the limit
@@
Let $\Gamma\in G_{g,n}$ be a graph, and let $\phi^{\Gamma}$ be a degree $d$ general numerical polarisation on $\Gamma$, compatible with $G_{g,n}$-morphisms. We define the classical $c$-stability condition $\mathfrak{m}^{\Gamma}(\phi^{\Gamma})$ induced by $\phi^{\Gamma}$ on $\Gamma$ by the assignment
$\mathfrak{m}^{\Gamma}(\phi^{\Gamma}):=\{\mathfrak{m}_{(e,h,A)}^{\Gamma}(\phi^{\Gamma}):=\Bigl{\lceil}\phi^{\Gamma}_{Z(e,h,A)}-\frac{e}{2}\Bigr{\rceil}:(e,h,A)\in\mathcal{D}(\Gamma)\},$
@@
The induced stability condition $\mathfrak{m}^{\Gamma}(\phi^{\Gamma})$ is the image, via the map of Proposition 2.12, of the $V$-stability condition
$\mathfrak{n}^{\Gamma}(\phi^{\Gamma}):=\{\mathfrak{n}^{\Gamma}(\phi^{\Gamma})_{Z}=\Bigl{\lceil}\phi^{\Gamma}_{Z}-\frac{\textup{val}(Z)}{2}\Bigr{\rceil}:Z\in\mathcal{C}(\Gamma)\},$
induced by $\phi^{\Gamma}$, defined in [Viv, Example 1.5].
Thus, to show that $\mathfrak{m}^{\Gamma}(\phi^{\Gamma})$ is well defined, it is sufficient to notice that whenever $\phi^{\Gamma}$ is compatible with $G_{g,n}$-morphisms, so is the induced $V$-stability condition. Indeed, this is true by the following observation. Let $\Gamma^{\prime}\in G_{g,n}$ and $c,\tilde{c}:\Gamma\to\Gamma^{\prime}$ be morphisms. For each $Z\in V(\Gamma^{\prime})$,
@@
A more explicit form for the classical stability condition $\mathfrak{m}^{\Gamma}(\phi^{\Gamma})$ will prove useful for constructing examples in the following section, thus we show:
@@
Let $g\geq 2$, and let $\Gamma\in G_{g,n}$ be a $2$-connected graph (i.e. a graph without separating edges). Let $\phi^{\Gamma}$ be a general degree $d$ numerical polarisation on $\Gamma$, compatible with $G_{g,n}$-morphisms. Then, the $\phi^{\Gamma}$-induced classical $c$-stability condition $\mathfrak{m}^{\Gamma}(\phi^{\Gamma})$ satisfies
$\mathfrak{m}_{(e,h,A)}^{\Gamma}(\phi^{\Gamma})=\Bigl{\lceil}\frac{(e+2h-2)(d+1-g)+(2g-e-2h)\phi^{\Gamma}_{A}-(e+2h-2)\phi^{\Gamma}_{A^{\mathsf{c}}}}{2g-2}\Bigr{\rceil}+h-1,$
where $\phi^{\Gamma}_{A}=\sum_{v:A(v)\subset A}\phi^{\Gamma}_{v}$, for $v\in V(\Gamma)$.
@@
We observe that the $V$-stability condition $\mathfrak{n}(\phi^{\Gamma})$ induced by a numerical polarisation $\phi^{\Gamma}$ of [Viv, Example 1.5] depends on the stability polytope containing $\phi^{\Gamma}$, rather than on the specific $\phi^{\Gamma}$.
Let $c:\Gamma^{\prime}\to\Gamma$, be a morphism, and let $\phi^{\Gamma}$ be a numerical polarisation on $\Gamma$. By generality of $\phi^{\Gamma}$, there exists a numerical polarisation $\phi^{\Gamma^{\prime}}$ on $\Gamma^{\prime}$ such that the pushforward $c(\phi^{\Gamma^{\prime}})$ is in the same polytope as $\phi^{\Gamma}$. Therefore it is enough to prove our statement on the initial objects of $G_{g,n}$, that is, trivalent graphs.
So, we assume that $\Gamma$ is trivalent. By compatibility with $G_{g,n}$-morphisms, since $\Gamma$ is $2$-connected, we have $\phi^{\Gamma}_{v_{1}}=\phi^{\Gamma}_{v_{2}}$ for any pair of unlabeled genus 00 vertices $v_{1},v_{2}$. Hence, the classical $c$-stability condition induced by $\phi^{\Gamma}$ on a $2$-connected trivalent graph in $G_{g,n}$ has elements
$\mathfrak{m}_{(e,h,A)}^{\Gamma}(\phi^{\Gamma})=\Bigl{\lceil}\frac{k(d+1-g)+(2g-2-k)\phi^{\Gamma}_{A}-k\phi^{\Gamma}_{A^{\mathsf{c}}}}{2g-2}\Bigr{\rceil}+h-1,$
@@
Let $\phi=\{\phi^{\Gamma}\}_{\Gamma\in G_{g,n}}$ be a collection of a $G_{g,n}$-morphisms compatible numerical polarisations $\phi^{\Gamma}$ for each $\Gamma$ in $G_{g,n}$. If, for each pair of graphs $\Gamma,\Gamma^{\prime}$, we have
@@
for any $(e,h,A)\in\mathcal{D}(\Gamma)\cap\mathcal{D}(\Gamma^{\prime})$, we define the $\phi$-induced semi-classical universal stability condition $\mathfrak{m}(\phi)$ as
@@
If, furthermore, $\phi$ is a universal numerical polarisation, then $\mathfrak{m}(\phi)$ is said to be classical.
@@
Clearly, not only each element $\phi^{\Gamma}$ of a contractions and automorphisms compatible collection $\{\phi^{\Gamma}\}_{\Gamma\in G_{g,n}}$ is compatible with $G_{g,n}$-morphisms, but their collection induces a family $\{\mathfrak{m}^{\Gamma}\}_{\Gamma\in G_{g,n}}$, whose integral data $\mathfrak{m}_{(e,h,A)}^{\Gamma}$ is the same for every $\Gamma$ such that $(e,h,A)\in\mathcal{D}(\Gamma)$. As a consequence, we obtain that every classical universal stability condition is semi-classical. The converse, however, does not hold, as shown by [PT23, Example 6.15] (see also Example 3.12 below).
The following lemma allows us to check whether a $c$-stability condition on a graph $\Gamma$ is classical by checking it only on half of the elements of $\mathcal{D}(\Gamma)$.
@@
Let $\mathfrak{m}^{\Gamma}$ be a degree $d$ classical $c$-stability condition on a $2$-connected graph $\Gamma\in G_{g,n}$, with $g\geq 2$, and let $(e,h,A)\in\mathcal{D}(\Gamma)$. Let $\phi^{\Gamma}\in\mathbb{R}^{n}$ be a numerical polarisation on $\Gamma$ such that $\mathfrak{m}_{(e,h,A)}^{\Gamma}=\mathfrak{m}_{(e,h,A)}^{\Gamma}(\phi^{\Gamma})$. Then $\mathfrak{m}^{\Gamma}_{(e,g+1-e-h,A^{\mathsf{c}})}=\mathfrak{m}^{\Gamma}_{(e,g+1-e-h,A^{\mathsf{c}})}(\phi^{\Gamma})$.